OpenStudy (anonymous):

solve using elimination -3x= 14+2y3y - 8x = -21

OpenStudy (mathstudent55):

-3x= 14+2y 3y - 8x = -21 First, subtract the 2y in the first equation to the left side. Write both equations with the x-term first, followed by the y-term, equaling a number. You should end up with this: -3x - 2y = 14 -8x + 3y = -21 Since you have -2y in the first equation, and 3y in the second equation, they don't add up to zero. You need to multiply the entire first equation by 3 and the entire second equation by 2. Then add up the two equations. The y-terms will be eliminated, and you can solve for x. Then substitute the value of x into one of the two original equations, and solve for y.